OpenStudy (anonymous):

Are 3x+33 _______ x+11 and 3 equivalent? Using complete sentences explain why or why not. Are 3x+33 _______ x+11 and 3 equivalent? Using complete sentences explain why or why not. @Mathematics

OpenStudy (anonymous):

Not equivalent since x=-11 returns no value for the fraction.

OpenStudy (anonymous):

0/0 is not 3

OpenStudy (agreene):

These are equivalent because if you do the division... \[\frac{3x+33}{x+11}=3\]

OpenStudy (anonymous):

They are equivalent for all x not equal to -11

OpenStudy (agreene):

no.. \[\frac{3(x+11)}{x+11}\] x+11 cancel \[=3\] these are the same thing.

OpenStudy (anonymous):

Plug in x=-11

OpenStudy (agreene):

I'm aware the fraction isnt defined at that point. However the fraction is also improper. you need to simplify before plugging in values.. and you will find that this is a horizontal line where y=3

OpenStudy (anonymous):

I know, I'm just pedantic.

OpenStudy (agreene):

a wrong pedantic.

OpenStudy (anonymous):

No. Equilvalence means it's correct across all of the domain.

OpenStudy (agreene):

It is. You have to reduce improper functions before evaluating them.